OpenStudy (anonymous):

help please solve the compound inequality 2x-5/4 <= - 6 or 2x-5/4 > 5

OpenStudy (dumbcow):

ok solve each inequality independently

OpenStudy (dumbcow):

2x-5/4 < -6 add 5/4 to both sides 2x < -6 + 5/4 combine fractions using common denominator 2x< -24/4 +5/4 2x< -19/4 divide by 2 on both sides x < -19/8 2x-5/4 > 5 same steps 2x > 20/4 + 5/4 2x> 25/4 x > 25/8 solution is x<=-19/8 and x>25/8

OpenStudy (anonymous):

thank you very much I did the 1st inequality wrong I was close but still had the wrong ans.

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

hey dumbcow the denominators are wrong should it be -19/2 and 25/2 at least that is what the system says

OpenStudy (dumbcow):

oh ok i misread the problem, please use parenthesis (2x-5)/4 <= -6 and (2x-5)/4 > 5

OpenStudy (dumbcow):

(2x-5)/4 <= -6 multiply by 4 2x-5 <= -24 add 5 2x <= -19 divide 2 x <= -19/2

OpenStudy (anonymous):

dumbcow could I get your help on this one also 2x-3/4 <= - 9 or 2x-3/4 > 4

OpenStudy (dumbcow):

is it (2x-3)/4 ??

OpenStudy (anonymous):

thats correct

OpenStudy (dumbcow):

ok its a big difference, please be careful when writing your questions so you get the correct answers anyway same steps Multiply by 4 Add 3 Divide by 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok i'm going to bed I'm tired of this for tonight thanks

OpenStudy (dumbcow):

haha ok you should get x<= -33/2 and x>19/2

OpenStudy (dumbcow):

night

OpenStudy (anonymous):

seriously I been at this stuff all day and it has given me hell cause I am not understanding all of it so I think its time to step away for a while.